OpenStudy (anonymous):

what are the solutions of the equation x^2-1=0

OpenStudy (jdoe0001):

\(\bf x^2-1=0\implies x^2=1\implies x=\sqrt{1}\)

OpenStudy (anonymous):

(x+1)(x-1)=0 (x+1)=0 (x-1)=0 x=-1 x=1

OpenStudy (jdoe0001):

well... \(\bf x^2-1=0\implies x^2=1\implies x=\pm \sqrt{1}\)